Suppose that the length of a rectangle is 3 inches longer than the width and that the perimeter of the rectangle is 78.
a. Set up an equation involving only W, the width of the rectangle.
b. Solve this equation algebraically to find the length of the rectangle. Find the width as well. Use the appropriate units of measure in you answers.
Answer: Length ______, Width ______

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Suppose that the length of a rectangle is 3 inches longer than the width and that the perimeter of the rectangle is 78.
a. Set up an equation involving only W, the width of the rectangle.
Perimeter = 2(W + L)
Perimeter = 2(W + (W+3)) = 2(2W+3) = 4W + 6
----------------------------------------------------

b. Solve this equation algebraically to find the length of the rectangle.
78 = 4W+6
4W = 72
W = 18 inches (Width)
---
L = 18 + 3 = 21 inches (Length)
-------------------------
Cheers,
Stan H.


Answer by solver91311(24713)



The perimeter of a rectangle is given by:



But for this rectangle we know that the length is 3 inches longer than the width, so:

which means that we can write:



and since we know that , we can write:



Solve for w to get the width, add 3 to get the length.


John